OpenStudy (anonymous):

Fred's diner serves ¾ pound steaks every night. Saturday night they served 16 steaks. How many pounds of steak did they serve in total?

OpenStudy (anonymous):

16 steaks * 3/4 pound/steak=48/4 pounds=12 lbs of steak